OpenStudy (anonymous):

A gas station stores its gasoline in a tank underground. The tank is a cylinder lying horizontally on its side. The radius is 4 ft, the length is 11 ft, and the top of the tank is 10 feet under the ground. The density of gasoline is 42 lb/ ft^3 a) Consider a slice of gasoline that is Δy ft thick and located y ft above the center of the cylinder. Write an expression giving the work required to pump the slice out.

OpenStudy (koikkara):

http://www.math.uri.edu/~eaton/mth142sp02hk8b.pdf is ur answer....

OpenStudy (koikkara):

shud i download and paste for u......

OpenStudy (koikkara):

click and save file...^^...??

OpenStudy (anonymous):

that's not the answer

OpenStudy (anonymous):

Consider a slice of gasoline that is Δy ft thick and located y ft above the center of the cylinder. Write an expression giving the work required to pump the slice out.

OpenStudy (anonymous):

@Koikkara can you help me or not?

OpenStudy (anonymous):

Okay so \[ W = F\cdot d \]

OpenStudy (anonymous):

We know that \(d =y\)

OpenStudy (anonymous):

\(F = mg = \Delta y A g = \Delta y \pi r^2g\) We know \(g\) is just acceleration due to gravity, and \(r\) is the radius of the tank.

OpenStudy (anonymous):

So: \[ W = \Delta y\pi r^2 gy \]

OpenStudy (anonymous):

\(m=\Delta y A\), \(A = \pi r^2\)

OpenStudy (anonymous):

Whoops... \(m = \Delta y A \delta\) where \(\delta\) is the density